OpenStudy (anonymous):

If f(x) = x^-(1/3) find the derivative of f(x)

OpenStudy (anonymous):

\[dx^{n}/dx=nx^{n-1}\] unless n=-1

OpenStudy (anonymous):

therefore \[dx^{-1/3}/dx=-1/3x^{-4/3}\]

OpenStudy (anonymous):

Or \[f(x) = x^{-1/3}=1/\sqrt[3]{x}\] \[f^{\prime}(x) =-1/3 \times x^{-4/3} = -1/(3x \sqrt[3]{x})\]

OpenStudy (anonymous):

wait is teh second way teh same asnwer as teh first?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

your looking for the \(lim_{h->0}\cfrac{f(x+h)-f(x)}{h}\) version tho right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

wait its ok....will will give u the same asnwer cuz the limh-> 0 is longer

OpenStudy (amistre64):

right

OpenStudy (anonymous):

i mean it wanst bad